\begin{thebibliography}{ZZZZ}

\bibitem[AnBi]{AnBi} B. Anchouche and I. Biswas, Einstein--Hermitian
connections on polystable principal bundles over a compact
K\"ahler manifold, \textit{Amer. Jour. Math.} \textbf{123}
(2001), 207--228.

\bibitem[AzBi]{AzBi} H. Azad and I. Biswas, On holomorphic principal bundles
over a compact Riemann surface admitting a flat connection, II, {\it Bull.
London Math. Soc.} {\bf 35}, (2003), 440--444.

\bibitem[At1]{At1} M. F. Atiyah, Vector bundles over an elliptic curve, {\it Proc.
London Math. Soc.} {\bf 7} (1957), 414--452.

\bibitem[At2]{At} M. F. Atiyah, Complex analytic connections in fibre
bundles, \textit{Trans. Amer. Math. Soc.} \textbf{85} (1957), 181--207.

\bibitem[FGN]{FGN} E. Franco, O. Garc{\'\i}a-Prada and P. E. Newstead, Higgs
bundles over elliptic curves, \textit{Ill. Jour. Math.} (in press).

\bibitem[FrMo]{FM} R. Friedman and  J. W. Morgan, Holomorphic principal bundles
over elliptic curves. II. The parabolic construction, {\it Jour. Diff. Geom.}
{\bf 56} (2000), 301--379. 

\bibitem[FMW]{FMW} R. Friedman, J. W. Morgan and E. Witten, Principal $G$-bundles
over elliptic curves, {\it Math. Res. Lett.} {\bf 5} (1998), 97--118. 

\bibitem[GrHa]{GH} P. Griffiths and J. Harris, \textit{Principles of algebraic geometry},
Reprint of the 1978 original, Wiley Classics Library. John Wiley \& Sons, Inc.,
New York, 1994.

\bibitem[La]{La} Y. Laszlo, About $G$-bundles over elliptic curves, {\it Ann. Inst.
Fourier} {\bf 48} (1998), 413--424.

\bibitem[Mu]{Mu} D. Mumford, \textit{Abelian varieties},
with appendices by C. P. Ramanujam and Yuri Manin, corrected reprint of the second
(1974) edition. Tata Institute of Fundamental Research Studies in Mathematics, 5,
Hindustan Book Agency, New Delhi, 2008. 

\bibitem[Ra]{Ra} A. Ramanathan, Stable
principal bundles on a compact Riemann surface, \textit{Math. Ann.}
\textbf{213} (1975), 129--152.

\bibitem[RaSu]{RS} A. Ramanathan and S. Subramanian, Einstein--Hermitian
connections on principal bundles and stability, \textit{Jour. Reine Angew.
Math.} \textbf{390} (1988), 21--31.

\bibitem[Si]{Si} C. T. Simpson, Moduli of representations of the fundamental group
of a smooth projective variety. II, {\it Inst. Hautes \'Etudes Sci. Publ. Math.}
{\bf 80} (1994), 5--79.

\bibitem[Th]{Th} M. Thaddeus, Mirror symmetry, Langlands duality, and commuting
elements of Lie groups, {\it Internat. Math. Res. Not.} (2001), 1169--1193.

\bibitem[We]{We} A. Weil, Generalisation des fonctions abeliennes,
\textit{Jour. Math. Pure Appl.} \textbf{17} (1938), 47--87.

\end{thebibliography}

\documentclass[12pt,reqno]{amsart}
\usepackage{amssymb}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}

\newtheorem{question}[theorem]{Question}

\numberwithin{equation}{section}

\begin{document}

\baselineskip=15pt

\title[Moduli space of $G$--connections on an elliptic curve]{Moduli space of
$G$--connections on an elliptic curve}

\author[I. Biswas]{Indranil Biswas}

\address{School of Mathematics, Tata Institute of Fundamental
Research, Homi Bhabha Road, Bombay 400005, India}

\email{indranil@math.tifr.res.in}

\subjclass[2000]{14D20, 14H52, 53B15}

\keywords{Elliptic curve, $G$-connection, moduli space, Riemann-Hilbert correspondence}

\thanks{The author is supported by J. C. Bose Fellowship.}

\date{}

\begin{abstract}
Let $X$ be a smooth complex elliptic curve and $G$ a connected reductive affine
algebraic group defined over $\mathbb C$. Let ${\mathcal M}_X(G)$ denote the moduli
space of topologically trivial algebraic $G$--connections on
$X$, that is, pairs of the form $(E_G\, , D)$, where $E_G$ is a topologically trivial
algebraic principal $G$--bundle on $X$, and $D$ is an algebraic connection on $E_G$.
We prove that ${\mathcal M}_X(G)$ does not admit any nonconstant algebraic function
while being biholomorphic to an affine variety.
\end{abstract}

\maketitle

\section{Introduction}

Take an irreducible smooth complex projective curve $X$ of genus one and also
take a connected
reductive complex affine algebraic group $G$. The moduli space of topologically trivial
algebraic principal $G$--bundles on $X$ is extensive studied (see \cite{At1},
\cite{FM}, \cite{FMW}, \cite{La}, \cite{FGN} and references therein). In particular,
an explicit description of this moduli space was obtained. Our aim here is to study the
moduli space of algebraic $G$--connections on $X$. More precisely, let
${\mathcal M}_X(G)$ be the moduli space pairs of the form $(E_G\, ,D)$, where
\begin{itemize}
\item $E_G$ is an algebraic principal $G$--bundle on $X$ such that the underlying
topological principal $G$--bundle is trivial, and

\item $D$ is an algebraic connection on $E_G$.
\end{itemize}
(The definition of an algebraic connection on $E_G$ is recalled in Section
\ref{se2.2}.)
The Riemann--Hilbert correspondence, which sends any flat connection on $X$ to its
monodromy representation, produces a biholomorphism from ${\mathcal M}_X(G)$ to the
Betti moduli space $M^B_X(G)$ that parametrizes all equivalence classes of
homomorphisms from $\pi_1(X)$ to $G$ that lie in the connected component of the trivial
homomorphism (see Section \ref{se4.1}). The moduli space $M^B_X(G)$ is an affine
variety. Since ${\mathcal M}_X(G)$ is biholomorphic to $M^B_X(G)$, we conclude that
there are many nonconstant holomorphic functions on ${\mathcal M}_X(G)$. This contrasts
with the following theorem proved here (see Theorem \ref{thm1}):

\begin{theorem}\label{thm0}
The variety ${\mathcal M}_X(G)$ does not admit any nonconstant algebraic function. 
\end{theorem}

Let $C$ be an irreducible smooth complex projective curve and
${\mathcal M}_C(G)$ the moduli space of $G$--connections on $C$ such that
the underlying topological principal $G$--bundle is trivial. As before,
${\mathcal M}_C(G)$ is biholomorphic to an affine variety, namely the
Betti moduli space. Theorem \ref{thm0} raises the following question:

\begin{question}
Does ${\mathcal M}_C(G)$ admit any nonconstant algebraic function?
\end{question}

\section{Algebraic connections on bundles}

\subsection{Connection on vector bundles}

Let $X$ be an irreducible smooth complex projective
curve of genus one. The holomorphic 
cotangent bundle of $X$ will be denoted by $K_X$. We note that $K_X$ is
algebraically trivial.

An \textit{algebraic connection} on an algebraic vector bundle 
$E\,\longrightarrow\, X$ is a holomorphic differential operator
$$
D\, :\, E\, \longrightarrow\, E\otimes K_X
$$
of order one satisfying the Leibniz rule which says that for any locally defined
holomorphic section $s$ of $E$ and any locally defined holomorphic function $f$ on $X$,
\begin{equation}\label{e1}
D(f\cdot s) \,=\, fD(s) + s\otimes df\, .
\end{equation}
Let $\overline{\partial}_E\, :\, E\, \longrightarrow\, E\otimes \Omega^{0,1}$ be 
the Dolbeault operator defining the holomorphic structure on $E$. It is 
straightforward to check that $D$ is an algebraic connection on $E$ if and only if 
$D+\overline{\partial}_E$ is a flat connection on $E$ such that the locally
defined (in analytic topology) flat sections of $E$ are holomorphic.

A rank $r$ {\it connection} on $X$ is a pair $(E\, , D)$, where 
$E$ is an algebraic vector bundle on $X$ of rank $r$, and $D$ is an algebraic 
connection on $E$.

An algebraic vector bundle $V$ on $X$ is called \textit{semistable} if
$$
\frac{\text{degree}(F)}{\text{rank}(F)}\,\leq\, \frac{\text{degree}(V)}{\text{rank}(V)}
$$
for all algebraic subbundles $F\,\subset\, V$ of positive rank.

\begin{lemma}\label{lem1}
Let $(E\, , D)$ be a rank $r$ connection on $X$. Then the 
algebraic vector bundle $E$ is semistable of degree zero.
\end{lemma}

\begin{proof}
Since $E$ admits a flat connection it follows immediately that $\text{degree}(E)\,=\,0$.

Assume that $E$ is not semistable. Let
$$
0\, \not=\, F\,\subsetneq\, E
$$
be the first term of the Harder--Narasimhan filtration of $E$, so the algebraic 
vector bundle $F$ is semistable, and
\begin{equation}\label{e2}
\frac{\text{degree}(F)}{\text{rank}(F)}\, > \frac{\text{degree}(V)}{\text{rank}(V)}
\end{equation}
for every algebraic subbundle $V\,\subset\, E/F$ of positive rank. Now consider 
the second fundamental form $S_D(F)$ for $F$, which is the following composition:
$$
F\,\hookrightarrow\, E\, \stackrel{D}{\longrightarrow}\, E\otimes K_X \,
\longrightarrow\, (E/F)\otimes K_X\, .
$$
{}From \eqref{e1} it follows immediately that $S_D(F)(fs)\,=\, fS_D(F)(s)$, where 
$s$ is any locally defined holomorphic section of $F$ and $f$ is any locally 
defined holomorphic function on $X$. Therefore, $S_D(F)$ is a holomorphic 
homomorphism of vector bundles. Since $K_X$ is trivial, from \eqref{e2} it follows 
that there is no nonzero holomorphic homomorphism of vector bundles from $F$ to 
$E/F$. So we conclude that $S_D(F)\,=\, 0$. Therefore, $D$ induces an algebraic 
connection on $F$. Since $F$ admits an algebraic connection, we have
$$
\text{degree}(F)\, =\, 0\, .
$$
On the other hand, since $\text{degree}(E)\,=\,0$, from \eqref{e2} it follows that
$$
\text{degree}(F)\, > \, 0\, .
$$
In view of the above contradiction, we conclude that $E$ is semistable.
\end{proof}

\begin{lemma}\label{lem2}
Let $E$ be a semistable vector bundle on $X$ of rank $r$ and degree zero.
Then $E$ admits an algebraic connection.
\end{lemma}

\begin{proof}
A theorem due to Atiyah and Weil says that an algebraic vector bundle $W$ over an
irreducible smooth complex projective curve
admits an algebraic connection if and only if 
the degree of any direct summand of $W$ is zero \cite{We}, \cite[p. 202, 
Proposition 17]{At}. Since $E$ is semistable of degree zero, every direct summand 
of $E$ is clearly of degree zero. Therefore, $E$ admits an algebraic connection.
\end{proof}

\subsection{Criterion for principal bundles}\label{se2.2}

Let $G$ be a connected reductive affine algebraic group defined over $\mathbb C$.
The center of $G$ is denoted by $Z(G)$. Let
$$
Z_0(G)\, \subset\, Z(G)
$$
be the connected component containing the identity element.
The Lie algebra of $G$ will be denoted by $\mathfrak g$.

An algebraic principal $G$--bundle $E_G$ over the elliptic curve $X$ is called
\textit{semistable} if for every triple of the form $(P\, ,E_P\, ,\chi)$, where
\begin{itemize}
\item $P\, \subset\, G$ is a parabolic subgroup,

\item $E_P\, \subset\, E_G$ is an algebraic reduction of structure group of $E_G$
to $P$, and

\item $\chi\, :\, P\, \longrightarrow\, {\mathbb C}^*$ is a character such that
$\chi\vert_{Z_0(G)}$ is trivial and the line bundle on $G/P$ associated to $\chi$
is ample,
\end{itemize}
the inequality $\text{degree}(E_P(\chi))\, \geq\, 0$, where $E_P(\chi)\,\longrightarrow
\, X$ is the algebraic line bundle associated to $E_P$ for the character $\chi$.
(See \cite{Ra}, \cite{RS}, \cite{AnBi}.)

For $G\, =\, \text{GL}(r,{\mathbb C})$, the above definition of semistability of $E_G$
is equivalent to the definition of semistability of the vector bundle of rank $r$
associated to $E_G$.

Given an algebraic principal $G$--bundle $p\, :\, E_G\, \longrightarrow\, X$, we have
the Atiyah exact sequence on $X$
$$
0\,\longrightarrow\, \text{ad}(E_G)\,:=\,E_G\times^G{\mathfrak g}
\,\longrightarrow\, \text{At}(E_G)\,:=\,(p_*TE_G)^G \,\longrightarrow\, TX
\,\longrightarrow\,0
$$
(see \cite{At}). An \textit{algebraic connection} on $E_G$ is a holomorphic
(hence algebraic) splitting of the above Atiyah exact sequence.

\begin{proposition}\label{prop1}
An algebraic principal $G$--bundle $E_G$ over $X$ admits an algebraic
connection if and only if the following two conditions hold:
\begin{enumerate}
\item $E_G$ is semistable, and

\item for every character $\chi$ of $G$, the degree of the associated line bundle
$E_G(\chi)\,\longrightarrow\, X$ is zero.
\end{enumerate}
\end{proposition}

\begin{proof}
We first recall a criterion for $E_G$ to admit an algebraic connection. The
principal $G$--bundle $E_G$ admits an algebraic
connection if and only if the following two conditions hold:
\begin{enumerate}
\item The adjoint vector bundle $\text{ad}(E_G)$
admits an algebraic connection, and

\item for every character $\chi$ of $G$, the degree of the associated line bundle
$E_G(\chi)\,\longrightarrow\, X$ is zero.
\end{enumerate}
(See \cite[p. 444, Theorem 3.1]{AzBi}.) On the other hand, the principal $G$--bundle
$E_G$ is semistable if and only the vector bundle $\text{ad}(E_G)$ is semistable
\cite[p. 214, Proposition 2.10]{AnBi}. Since any algebraic connection on $E_G$
induces an algebraic connection on $\text{ad}(E_G)$, the proposition follows by
combining these with Lemma \ref{lem1} and Lemma \ref{lem2}.
\end{proof}

\section{Moduli space of connections}

Let ${\mathcal M}_X$ denote the moduli space of rank one algebraic connections
on $X$, so ${\mathcal M}_X$ parametrizes isomorphism classes of pairs of
the form $(L\, ,D)$, where $L$ is an algebraic line bundle on $X$ of degree zero,
and $D$ is an algebraic connection on $L$. Let
$$
J \,=\, J(X)\,=\, \text{Pic}^0(X)
$$
be the Jacobian of $X$. Once we fix a point $x_0\,\in\, X$, there is the
isomorphism $X\, \longrightarrow\, J$ defined by $x\,\longmapsto\, {\mathcal O}_X
(x-x_0)$. Let
\begin{equation}\label{e3}
\varphi\, :\, {\mathcal M}_X\,\longrightarrow\, J
\end{equation}
be the projection defined by $(L\, ,D)\,\longmapsto\, L$. This $\varphi$ is
a smooth surjective algebraic morphism. The space of all algebraic connections
on an algebraic line bundle $L$ is an affine space for $H^0(X,\, K_X)$.
Therefore, $\varphi$ in \eqref{e3} makes
${\mathcal M}_X$ a torsor over $J$ for the trivial vector bundle $J\times
H^0(X,\, K_X)$. Equivalently, ${\mathcal M}_X$ is an algebraic principal bundle
over $J$ with structure group $H^0(X,\, K_X)$.

Using Serre duality, $H^0(X,\, K_X)$ is identified with $\mathbb C$. Therefore,
${\mathcal M}_X$ is an algebraic principal $\mathbb C$--bundle over $J$.

\begin{lemma}\label{lem3}
The algebraic principal $\mathbb C$--bundle ${\mathcal M}_X\,
\stackrel{\varphi}{\longrightarrow}\, J$ is nontrivial.
\end{lemma}

\begin{proof}
Isomorphism classes of algebraic principal $\mathbb C$--bundles
on $J$ are parametrized by $H^1(J,\, {\mathcal O}_J)$.
Indeed, given a principal $\mathbb C$--bundle $E$ on $J$, choosing local
trivializations (in analytic topology) of $E$ and taking their differences, the
cohomology class in $H^1(J,\, {\mathcal O}_J)$ corresponding to $E$
is constructed. We will now recall the Dolbeault analog of this construction.

Let $\beta\, :\, E\,\longrightarrow\, J$ be an algebraic principal
$\mathbb C$--bundle on $J$. Since $\mathbb C$ is contractible, the
fiber bundle $E$ admits a $C^\infty$ section. Let
$$
\gamma\, :\, J\, \longrightarrow\, E
$$
be a $C^\infty$ section, meaning $\beta\circ\gamma\,=\, \text{Id}_J$. We note
that $\gamma$ need not be holomorphic; in fact, $\beta$ admits a holomorphic
section if and only if the algebraic principal $\mathbb C$--bundle $E$
is trivial. Take a point $x\,\in\, J$. Let
$$
d\gamma (x)\, :\, T^{\mathbb R}_xJ\, \longrightarrow\, T^{\mathbb R}_{\gamma(x)} E
$$
be the differential of $\gamma$ at $x$, where $T^{\mathbb R}$ denotes the
real tangent space. Let ${\mathbb J}_x$ (respectively, ${\mathbb J}_{\gamma(x)}$)
be the almost complex structure on $J$ (respectively, $E$) at $x$
(respectively, $\gamma(x)$). Now consider the homomorphism
$$
T^{\mathbb R}_xJ\, \longrightarrow\, T^{\mathbb R}_{\gamma(x)} E\, , ~\
v\,\longmapsto\, d\gamma (x)({\mathbb J}_x(v))-
{\mathbb J}_{\gamma(x)}(d\gamma (x)(v))\, .
$$
Since the map $\beta$ is holomorphic, this homomorphism produces a
homomorphism
$$
\omega(x)\, :\, T^{0,1}_xJ\, \longrightarrow\, \mathbb C\, ;
$$
using the action of $\mathbb C$ on $E$, the vertical tangent bundle on $E$
for the projection $\beta$ is algebraically identified with the trivial line bundle
with fiber $\mathbb C$ (the Lie algebra of the group $\mathbb C$).
Therefore, we get a $(0\, ,1)$--form $\omega$ on $J$ whose evaluation
at any point $x\, \in\, J$ is $\omega(x)$ defined above. We note that $\omega$ is
the obstruction for the map $\gamma$ to be holomorphic. The element in
$H^1(J,\, {\mathcal O}_J)$ represented by $\omega$ is the cohomology class
corresponding to the algebraic principal $\mathbb C$--bundle $E$.

Now consider the algebraic principal $\mathbb C$--bundle ${\mathcal M}_X$ in
\eqref{e3}. Any algebraic line bundle on $X$ of degree zero admits a unique
unitary flat connection. Let
\begin{equation}\label{ga}
\gamma\, :\, J\, \longrightarrow\, {\mathcal M}_X
\end{equation}
be $C^\infty$ section defined by $L\, \longmapsto\, (L\, , D^u_L)$, where
$D^u_L$ is the unique unitary flat connection on $L$. Consider the
$(0\, ,1)$--form $\omega$ on $J$ constructed as above from this section $\gamma$.
It is straightforward to check that
\begin{itemize}
\item $\omega$ is invariant under the translations of the group $J$, and

\item $\omega$ is nonzero.
\end{itemize}
These two imply that the cohomology class in $H^1(J,\, {\mathcal O}_J)$
represented by $\omega$ is nonzero. This completes the proof.
\end{proof}

\begin{lemma}\label{lem4}
Let $E$ and $F$ be two algebraic principal $\mathbb C$--bundles on $J$ such that
both $E$ and $F$ are nontrivial. Then the total spaces of $E$ and $F$ are
algebraically isomorphic.
\end{lemma}

\begin{proof}
Let $\eta\, :\, E\times {\mathbb C}\,\longrightarrow\, E$ be
the action of $\mathbb C$ on the principal $\mathbb C$--bundle $E$.
Take any nonzero complex number $\lambda$. Let $E^\lambda$ be the
algebraic principal $\mathbb C$--bundle on $J$ whose total space is
$E$ but the action of any $c\, \in\, \mathbb C$ on $E$ is the morphism
$z\,\longmapsto\, \eta(z\, ,\lambda\cdot c)$. If $\omega_0\,\in\,
H^1(J,\, {\mathcal O}_J)$ is the cohomology class corresponding to $E$, then
the cohomology class corresponding to $E^\lambda$ is $\omega_0/\lambda$.
Since $\dim H^1(J,\, {\mathcal O}_J)\,=\, 1$, it now follows that the
total spaces of any two nontrivial algebraic principal $\mathbb C$--bundles on $J$ 
are algebraically isomorphic.
\end{proof}

Let
\begin{equation}\label{V}
0\,\longrightarrow\, {\mathcal O}_J \,\longrightarrow\, V\,
\stackrel{\sigma}{\longrightarrow}\, {\mathcal O}_J \,\longrightarrow\, 0
\end{equation}
be a nontrivial extension of ${\mathcal O}_J$ by ${\mathcal O}_J$; such an
extension exists because $H^1(J,\, {\mathcal O}_J)\,\not=\, 0$. Let
$$
s_1\,:\, J\, \longrightarrow\, {\mathcal O}_J
$$
be the section given by the constant function $1$ on $J$. Define
\begin{equation}\label{Z}
Z\,:=\, \sigma^{-1}(s_1(J))\,\subset\, V\, ,
\end{equation}
where $\sigma$ is the homomorphism in \eqref{V}. Using the inclusion
$$
{\mathcal O}_J \,\hookrightarrow\, Z\,\subset\, V
$$
in \eqref{V} it follows that $Z$ is an algebraic principal $\mathbb C$--bundle over $J$.

\begin{proposition}\label{prop2}
The variety $Z$ in \eqref{Z} is isomorphic to ${\mathcal M}_X$.
\end{proposition}

\begin{proof}
We will show that the algebraic principal $\mathbb C$--bundle $Z\,\longrightarrow\,
J$ is nontrivial. Since local trivializations of $Z$ give local splittings of
the exact sequence in \eqref{V}, the cohomology class in $H^1(J,\, {\mathcal O}_J)$
corresponding to the principal $\mathbb C$--bundle $Z$ coincides with the cohomology
class corresponding to the extension in \eqref{V}. Since the extension in \eqref{V}
is non-split, we conclude that the cohomology class in $H^1(J,\, {\mathcal O}_J)$
corresponding to the principal $\mathbb C$--bundle $Z$ is nonzero. Now the
proposition follows from Lemma \ref{lem3} and Lemma \ref{lem4}.
\end{proof}

\begin{proposition}\label{prop3}
The variety ${\mathcal M}_X$ does not admit any nonconstant algebraic function.
\end{proposition}

\begin{proof}
Let
\begin{equation}\label{P}
p\, :\,{\mathbb P}\,:=\, {\mathbb P}(V)\,\longrightarrow\, J
\end{equation}
be the projective bundle over $J$ associated to $V$ in \eqref{V}. Note that the
projection $\sigma$ in \eqref{V} defines a section of the projective bundle
${\mathbb P}$ in \eqref{P}. The image of this section will be denoted by $S$. We have
\begin{equation}\label{i}
{\mathbb P}\setminus S\,=\, Z\, ,
\end{equation}
where $Z$ is constructed in \eqref{Z}. Let ${\mathcal O}_{\mathbb P} (1)$ denote
the tautological quotient line bundle of
$p^*V$. For any $n\, \geq\,1$, the tensor product ${\mathcal O}_{\mathbb P}
(1)^{\otimes n}$ will be denoted by ${\mathcal O}_{\mathbb P} (n)$.

The restrictions of both ${\mathcal O}_{\mathbb P}(1)$ and ${\mathcal O}_{\mathbb P}(S)$
to any fiber of $p$ are of degree one. Therefore, from the see--saw theorem,
\cite[p. 51, Corollary 6]{Mu}, we know
that there is a line bundle $\xi$ on $J$ such that
\begin{equation}\label{P2}
{\mathcal O}_{\mathbb P} (1)\otimes p^*\xi\,=\, {\mathcal O}_{\mathbb P}(S)\, .
\end{equation}
The restriction of ${\mathcal O}_{\mathbb P} (1)$ to $S$ is the trivial line bundle
because the quotient line bundle in \eqref{V} is trivial. From the Poincar\'e
adjunction formula we know that the restriction of ${\mathcal O}_{\mathbb P}(S)$ to
$S$ is the normal bundle to $S$ (see \cite[p. 146]{GH} for
the Poincar\'e adjunction formula). The normal bundle to $S$ is clearly identified
with the restriction to $S$ of the relative tangent bundle for the projection $p$
in \eqref{P}. From \eqref{V} we know that the restriction to $S$ of the
relative tangent bundle for $p$ is $Hom({\mathcal O}_J\, ,{\mathcal O}_J)\,=\,
{\mathcal O}_J$. Since both ${\mathcal O}_{\mathbb P} (1)\vert_S$ and
${\mathcal O}_{\mathbb P}(S)\vert_S$ are trivial, from \eqref{P2} we conclude that
the line bundle $\xi$ is trivial. Hence we have
\begin{equation}\label{P3}
{\mathcal O}_{\mathbb P} (1)\,=\, {\mathcal O}_{\mathbb P}(S)\, .
\end{equation}
So, ${\mathcal O}_{\mathbb P}(n)\,=\, {\mathcal O}_{\mathbb P} (n\cdot S)$
for all $n\, \geq\, 1$. Therefore, the projection formula says that
$$
p_*{\mathcal O}_{\mathbb P}(n\cdot S)\,=\, p_* ({\mathcal O}_{\mathbb P} (n))\,=\,
\text{Sym}^n(V)\, .
$$
This implies that
\begin{equation}\label{e6}
H^0({\mathbb P},\, {\mathcal O}_{\mathbb P}(n\cdot S))\,=\,
H^0(J,\, f_*({\mathcal O}_{\mathbb P}(n\cdot S)))\,=\, H^0(J,\,\text{Sym}^n(V))\, .
\end{equation}
The vector bundle $\text{Sym}^n(V)$ is isomorphic to the vector bundle $F_{n+1}$
in \cite[p. 432, Theorem 5]{At1}. It is shown in \cite{At1} that
$$
\dim H^0(J,\, F_{n})\,=\, 1
$$
(this is proved in \cite[p. 430, Lemma 15]{At1}; see also the penultimate line
in \cite[p. 432]{At1}). Therefore, from \eqref{e6} we conclude that
$$
H^0({\mathbb P},\, {\mathcal O}_{\mathbb P}(n\cdot S))\,=\, \mathbb C\, \ \ \ \forall \
n \, >\, 0\, .
$$
Now from \eqref{i} it follows that
$$
H^0(Z,\, {\mathcal O}_{Z})\,=\, \mathbb C\, .
$$
In view of this, the proposition follows from Proposition \ref{prop2}.
\end{proof}

\section{Functions of moduli space of $G$--connections}

Let $G$ be a connected reductive affine algebraic group over $\mathbb C$.
Let ${\mathcal M}_X(G)$ denote the moduli space of topologically
trivial algebraic $G$--connections on
$X$, meaning pairs of the form $(E_G\, , D)$, where $E_G$ is an algebraic
principal $G$--bundle on $X$ such that $E_G$ is topologically trivial,
and $D$ is an algebraic connection on $E_G$.

\begin{theorem}\label{thm1}
The variety ${\mathcal M}_X(G)$ does not admit any nonconstant algebraic function.
\end{theorem}

\begin{proof}
We saw in Proposition \ref{prop3} that the variety ${\mathcal M}_X$
does not admit any nonconstant algebraic function. Therefore, for any positive
integer $d$, the Cartesian product $({\mathcal M}_X)^d$ does not admit any
nonconstant algebraic function.

Fix a maximal torus ${\mathcal T}\, \subset\, G$. Let ${\mathcal M}_X({\mathcal T})$
be the moduli space of ${\mathcal T}$--connections on $X$, meaning pairs of the
form $(E_{\mathcal T}\, ,D)$, where $E_{\mathcal T}$ is an algebraic principal
${\mathcal T}$--bundle on $X$ and $D$ is an algebraic connection on $E_{\mathcal T}$.
We note that since ${\mathcal T}$ is a product of copies of
the multiplicative group ${\mathbb C}^*\,=\, {\mathbb C}\setminus \{0\}$, if an
algebraic principal ${\mathcal T}$--bundle $E_{\mathcal T}$ admits an algebraic
connection, then $E_{\mathcal T}$ is topologically trivial. The inclusion of $\mathcal T$
in $G$ produces a morphism
$$
\psi\, :\, {\mathcal M}_X({\mathcal T})\,\longrightarrow\, {\mathcal M}_X(G)\, .
$$
This morphism $\psi$ is known to be surjective \cite[p. 1177, Theorem 4.1]{Th}.

Let $\delta$ denote the dimension of ${\mathcal T}$, so ${\mathcal T}$ is isomorphic
to $({\mathbb C}^*)^\delta$. Therefore, ${\mathcal M}_X({\mathcal T})$ is isomorphic to
the variety $({\mathcal M}_X)^\delta$. Since $({\mathcal M}_X)^\delta$ does not admit
any nonconstant algebraic function, and $\psi$ is surjective, we conclude that
${\mathcal M}_X(G)$ does not admit any nonconstant algebraic function.
\end{proof}

\subsection{The Betti moduli space}\label{se4.1}

We now recall the definition of the Betti moduli space $M^B_X(G)$ associated to the
pair $(X\, ,G)$ \cite{Si}. The identity element of $G$ will be denoted by $e$. We will
identify $\pi_1(X)$ with ${\mathbb Z}\bigoplus {\mathbb Z}$. Consider the morphism
$$
f\, :\, G\times G\,\longrightarrow\, G\, , ~\ (x\, ,y)\, \longmapsto\,
xyx^{-1}y^{-1}\, .
$$
Let ${\mathcal R}_X(G)$ denote the connected component, containing $(e\, ,e)$, of
$f^{-1}(e)$. It is an affine variety because $G$ is so. The simultaneous conjugation
action of $G$ on $G\times G$ preserves ${\mathcal R}_X(G)$. The geometric invariant
theoretic quotient ${\mathcal R}_X(G)/\!\!/G$ is the Betti moduli space $M^B_X(G)$.
It is an affine variety. The Riemann--Hilbert correspondence produces a biholomorphism
between ${\mathcal M}_X(G)$ and $M^B_X(G)$; it sends an algebraic connection to
the monodromy of the corresponding flat connection.

In contrast with Theorem \ref{thm1}, the variety ${\mathcal M}_X(G)$ has plenty
of holomorphic functions because it is biholomorphic to an affine variety.


\end{document}